OpenStudy (anonymous):

Find an equation of variation where y varies directly as x and y = 75 when x = 5, find y when x is 1.

OpenStudy (anonymous):

so when you hear y varies directly as x, that means y = k*x. put in 75 for y and 5 for x and solve for k.

OpenStudy (anonymous):

"where y varies directly as x"\[ y = kx \]"y = 75 when x = 5"\[ 75 = k(5) \]

OpenStudy (anonymous):

when you solve for k, you can put it in for y = k*1

OpenStudy (anonymous):

"find y when x is 1"\[ k(1) = ? \]

OpenStudy (anonymous):

\[y = k*x\] \[75 = k*5\] \[75/5 = k\] \[15 = k\] \[y = 15(x)\] \[y = 15*1\] y = 15 when x = 1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem, happy to help!